Income Statements under Absorption and Variable Costing

Shawnee Motors Inc. assembles and sells MP3 players. The company began operations on August 1 and operated at 100% of capacity during the first month. The following data summarize the results for August:

Sales (18,000 units) $2,880,000
Production costs (23,000 units):
Direct materials $1,361,600
Direct labor 653,200
Variable factory overhead 326,600
Fixed factory overhead 218,500 2,559,900
Selling and administrative expenses:
Variable selling and administrative expenses $396,800
Fixed selling and administrative expenses 153,600 550,400

If required, round interim per-unit calculations to the nearest cent.

a. Prepare an income statement according to the absorption costing concept.

Shawnee Motors Inc.
Absorption Costing Income Statement
For the Month Ended August 31
Sales $
Cost of goods sold
Gross profit $
Selling and administrative expenses
Income from operations $

Feedback

a. Under absorption costing, the cost of goods manufactured includes direct materials, direct labor, and factory overhead costs. Both fixed and variable factory costs are included as part of factory overhead.

Learning Objective 1.

b. Prepare an income statement according to the variable costing concept.

Shawnee Motors Inc.
Variable Costing Income Statement
For the Month Ended August 31
Sales $
Variable cost of goods sold
Manufacturing margin $
Variable selling and administrative expenses
Contribution margin $
Fixed costs:
Fixed factory overhead $
Fixed selling and administrative expenses
Total fixed costs
Income from operations $
Answer #1
a. Shawnee Motors Inc.
Absorption costing income statement
For the month ended August 31
Sales $ 28,80,000
Less: Cost of goods sold (18,000 X $111.30) $ 20,03,400
Gross profit $   8,76,600
Less: selling and administrative expenses $   5,50,400
Income from operations $   3,26,200
Workings:
Production cost per unit absorption costing:
Direct material per unit = $1361600 / 23,000 = $   59.20
Direct labor per unit = $653200 / 23,000 = $   28.40
Variable Factory overhead = $326600 / 23,000 = $   14.20
Fixed Factory overhead = $218500 / 23,000 = $      9.50
Total production cost per unit = = $ 111.30
b.
Shawnee Motors Inc.
Variable costing income statement
For the month ended August 31
Sales $                28,80,000
Less: Variable cost of goods sold (18,000 X $101.80) $                18,32,400
Manufacturing expenses $                10,47,600
Less: Variable selling and administrative expenses $                  3,96,800
Contribution margin $                  6,50,800
Less: Fixed costs:
Fixed factory overhead $   2,18,500
Fixed selling and administrative expenses $   1,53,600
Total fixed costs $                  3,72,100
Income from operations $                  2,78,700
Workings:
Production cost per unit variable costing:
Direct material per unit = $1361600 / 23,000 = $   59.20
Direct labor per unit = $653200 / 23,000 = $   28.40
Variable Factory overhead = $326600 / 23,000 = $   14.20
Total production cost per unit = $ 101.80

If it is helpful, please rate the answer and if any doubt arises let me know